# Inferring entropy production in many-body systems using nonequilibrium MaxEnt

### v3

@@ Abstract
This quantity obeys the bounds $0\leq{\Sigma}_{\bm{g}}\leq{\Sigma}_{\bm{g}}^{\mathrm{DPI}}$, as shown in the Supplemental Material (SM) [37]. ${\Sigma}_{\bm{g}}$ quantifies the irreversibility captured by observable expectations, and it vanishes if and only if the expectation of $\bm{g}$ is the same under the forward and reverse processes. Large-deviations theory gives ${\Sigma}_{\bm{g}}$ a physical interpretation in terms of fluctuations: given $n$ sample trajectories from the reverse process, the probability that the empirical average of observables $\bm{g}$ is equal to $\left\langle\bm{g}\right\rangle_{p}$ scales as $\asymp e^{-n{\Sigma}_{\bm{g}}}$ [38].
@@
So far, we have made no assumptions regarding multipartite structure. As we show in End Matter, in the presence of multipartite structure, where only a single observable can be active in a given trajectory, Eq. (3) can be split into smaller optimization problems, each involving a proportional fraction of variables and samples. This dramatically reduces the computation and memory required to solve Eq. (3) (see End Matter for an example).
Maximum likelihood EP decomposition.— As we show in the SM [37], the distribution $q^{*}$ that optimizes (2) belongs to an exponential family,
@@
The first term is the additional EP that could be inferred by constraining all statistics (not just the mean) of $\bm{g}$ in Eq. (2). The second term is the EP that cannot be inferred from any statistics of $\bm{g}$.
@@
${\Sigma}\geq{\Sigma}_{\bm{g}}^{\textrm{TUR}}\vcentcolon=\ln\left(1+2\langle\bm{g}\rangle_{p}^{\top}\mathsf{K}_{p}^{-1}\langle\bm{g}\rangle_{p}\right)$ (12)
@@
Our Pythagorean theorem (6) is related to previous decompositions of EP in interacting systems [53] and systems with nonconservative forces [54, 55, 35]. Moreover, the general form of Eq. (2) is a Maximum Entropy (MaxEnt) problem over trajectory distributions [56, 57, 58, 59], sometimes called “maximum caliber” (MaxCal) [60]. Related techniques have been used to infer models from which the entropy flow can be estimated, though without ensuring a lower bound on EP [61].
@@
To calculate ${\Sigma}_{\bm{g}}$, we sample from steady state using Monte Carlo. We then optimize Eq. (3) using gradient ascent with Barzilai–Borwein step sizes [71]. We perform early-stopping using held-out validation data, which avoids overfitting even in the far-from-equilibrium regime where many reverse transitions are not sampled. Reported EP estimates are computed on held-out test data. Details of the data generation, optimization, and analysis can be found in the SM [37] and our code repository [72].
@@ I Data Availability Statement
The code used in this study is available at a GitHub repository [72].

### v4

# Inferring entropy production in many-body systems using nonequilibrium maximum entropy

@@ Abstract
This quantity obeys the bounds $0\leq{\Sigma}_{\bm{g}}\leq{\Sigma}_{\bm{g}}^{\mathrm{DPI}}$, as shown in the Supplemental Material (SM) [37] ${\Sigma}_{\bm{g}}$ quantifies the irreversibility captured by observable expectations, and it vanishes if and only if the expectation of $\bm{g}$ is the same under the forward and reverse processes. Large-deviations theory gives ${\Sigma}_{\bm{g}}$ a physical interpretation in terms of fluctuations: given $n$ sample trajectories from the reverse process, the probability that the empirical average of observables $\bm{g}$ is equal to $\left\langle\bm{g}\right\rangle_{p}$ scales as $\asymp e^{-n{\Sigma}_{\bm{g}}}$ [41].
@@
So far, we have made no assumptions regarding multipartite structure. Here, we say that the observables are “multipartite” if they can be partitioned into blocks, such that only a single block can be active in any given trajectory. In practice, this situation often arises when observables depend on local subsystems and only one subsystem can change state at a given time. We provide a formal definition of multipartite observables in the End Matter. There, we show that multipartiteness allows Eq. (3) to be split into smaller optimization problems, each involving a proportional fraction of variables and samples, dramatically reducing the computation and memory required for optimization.
Maximum likelihood EP decomposition.—
As we show in the SM [37], the distribution $q^{*}$ that optimizes (2) belongs to an exponential family,
@@
is the additional EP that could be inferred by constraining all statistics (not just the mean) of $\bm{g}$ in Eq. (2). The second term is the EP that cannot be inferred from any statistics of $\bm{g}$.
@@
${\Sigma}\geq{\Sigma}_{\bm{g}}^{\textrm{TUR}}\vcentcolon=\ln\left(1+2\langle\bm{g}\rangle_{p}^{\top}\mathsf{K}_{p}^{-1}\langle\bm{g}\rangle_{p}\right).$ (12)
@@
Our Pythagorean theorem (6) is related to previous decompositions of EP in interacting systems [56] and systems with nonconservative forces [57, 35]. Moreover, the general form of Eq. (2) is a Maximum Entropy (MaxEnt) problem over trajectory distributions [58, 59, 60, 61], sometimes called “maximum caliber” (MaxCal) [62]. Related techniques have been used to infer models from which the entropy flow can be estimated, though without ensuring a lower bound on EP [63].
@@
To calculate ${\Sigma}_{\bm{g}}$, we sample from steady state using the Monte Carlo method. We then optimize Eq. (3) using gradient ascent with Barzilai–Borwein step sizes [73]. We perform early-stopping using held-out validation data, which avoids overfitting even in the far-from-equilibrium regime where many reverse transitions are not sampled. Reported EP estimates are computed on held-out test data. Details of the data generation, optimization, and analysis can be found in the SM [37] and our code repository [74].
@@ I Data Availability Statement
The code used in this study is available in a public repository [74].